# Mutual benefits of social learning and algorithmic mediation for cumulative culture

### v1

@@
Agnieszka Czaplicka1, Fabian Baumann1,2, Iyad Rahwan1
@@
2Department of Biology, University of Pennsylvania, Philadelphia, PA 19104
@@
The remarkable ecological success of humans is often attributed to our ability to develop complex cultural artefacts that enable us to cope with environmental challenges. The evolution of complex culture (cumulative cultural evolution) is usually modeled as a collective process in which individuals invent new artefacts (innovation) and copy information from others (social learning). This classic picture overlooks the fact that in the digital age, intelligent algorithms are increasingly mediating information between humans, with potential consequences for cumulative cultural evolution. Building on an established model of cultural evolution, we investigate the combined effects of network-based social learning and a simplistic version of algorithmic mediation on cultural accumulation. We find that algorithmic mediation has a strong impact on cultural accumulation and that this impact generally increases as social networks become less densely connected. Moreover, cultural accumulation tends to be optimal when social learning and algorithmic mediation are combined, and the optimal ratio depends on the network’s density. Our modeling results are a first step towards formalising the impact of intelligent algorithms on cumulative cultural evolution within an established framework. Models of this kind will also help to uncover mechanisms of human-machine interaction in cultural contexts, guiding hypotheses for future experimental testing.
@@ 1 Introduction
Humans’ ability to create complex cultural artefacts such as languages, scientific theories, art and technology, is often seen as one of the most important aspects of our success as a species [1, 2]. An important driving force behind this cultural accumulation is the ability of humans to learn socially, i.e. from their peers [3, 4], and pass our their acquired knowledge on to others [3, 5, 1]. Social learning is a highly adaptive process, influenced by cognitive biases that can manifest as different social learning strategies ranging from copying successful individuals (prestige bias) [6, 7], to frequent cultural artefacts (context bias) [8, 9] to specific information (content bias) [10]. At the population level, social learning is a necessary condition for cumulative cultural evolution, a process that describes the emergence of increasingly complex cultural artifacts through a ratchet effect: By preserving important information over time—i.e., from generation to generation—social learning allows collectives to produce cultural artefacts that lie outside of the scope of individuals [11, 12]. Yet the accumulation of culture in populations can also be limited by factors such as population size [12] or the structure of the underlying social network [13].
In the digital age, much of human social learning is shaped by intelligent algorithms. Instead of learning directly from their peers, individuals absorb information mediated by algorithms. Examples of such algorithms include recommendation systems that influence consumer behavior through personalized product information [14], algorithmically curated content feeds on social media platforms [15] and search engines that rank websites and online content [16]. By facilitating the targeted retrieval of information, these systems can significantly improve the ability of individuals to navigate complex online information landscapes and promote the transfer of knowledge within collectives. However, information-mediating algorithms also have a potential downside: algorithmic personalization has been shown to restrict the scope for individual exploration and limit the diversity of information, for example by actively creating filter bubbles [17] or exacerbating natural tendencies towards homophily [18, 19]. Recognizing these impacts, previous theoretical and empirical works have investigated the impact of algorithms on the diversity of content [20], the formation of opinions [21, 22], and the spread of misinformation [23]. However, little attention has been paid to the question of how algorithmic mediation may affect fundamental and more general processes of cultural evolution [24, 25]. In particular, it is unclear how algorithmic mediation interacts with human social learning to influence the accumulation of complex cultural artifacts.
@@
In this paper, we take a step forward to close this gap by building on an established model of cultural evolution [26]. More specifically, we extend the model by a simple algorithm that first aggregates all available information and then redistributes it in a personalized way. This allows us to study the interplay between social learning (i.e., people share information based on a social network) and algorithmic mediation (i.e., information is mediated by an algorithm). Our modeling results shed light on an emerging trade-off that humans face in the digital age: cultural accumulation is neither optimal through pure social learning nor through purely algorithmically mediated information transfer. Instead, both types of learning are most powerful when combined, where the optimal ratio depends on the underlying social network. The dynamic interplay between social learning and algorithmic mediation suggests that understanding cultural evolution in the digital age requires considering both types of learning together, rather than in isolation.
@@ 2 Model
We build on previous theoretical work [26] that examined how cultural accumulation is affected by different constraints on the individual level. In particular, while the impact of different (social) learning strategies and the costs associated to learning on cultural accumulation were studied, the influence of a social network structure and algorithmic mediation was not investigated.
To fill this gap, we first consider cultural accumulation in populations of agents that communicate based on random (social) networks with different densities of network connections. In a second step, we extend the model by a simple algorithm that mediates information among agents. The individual-based model presented in [26] considers a dynamic process that evolves in generations, with each ”time step” comprising the birth and death of an entire generation of agents. First, the agents copy from the previous generation and then, having learnt all possible information, they start to innovate. Here we deviate from this approach. We modify the model by re-scaling the dynamics so that at each time step a single cultural trait can be acquired either by copying or by innovation [27]. This modification allows us to study the interplay between different mechanisms of information transmission between agents—social learning and algorithmic mediation—and the consequences for population-scale cultural accumulation.
@@ 2.1 Cultural space
We consider a space of cultural traits (or cultural space) that defines how new traits (innovations) arise based on existing ones. In line with previous works, the cultural space consists of several cultural branches (or paths), and that traits are functionally dependent, reflecting the cumulative nature of human culture [13, 28, 26, 29].
More specifically, the cultural space is structured as follows. Each trait $x_{l}$ is defined by two components: the cultural branch $x$ it belongs to and its complexity level $l$, see Fig. 1A. Cultural branches $x$ represent the fact that cultural evolution often follows multiple parallel paths of development [28, 13], and levels $l$ model the functional dependency and increasing complexity of cultural artifacts built on top of each other. We consider a finite set of possible branches, i.e., $x=1,...,X$ and assume that cultural complexity levels $l$ are unbounded, i.e. $l=1,2,...,\infty$. The functional dependency of the traits $x$ additionally requires that the traits are acquired in ascending order, starting with the lowest level $l=1$, or more generally: an agent that wants to acquire a trait on level $l$ must possess a trait on level ${l-1}$.
Each cultural trait $x_{l}$ is characterised by an inherent quality score, or payoff, $z(x_{l})$, and we assume that the extent to which the traits are beneficial to agents increases with $z$ [5, 13, 30, 31]. Note, however, that we do not link an individual’s probability to survive to their payoff. Instead, cultural selection happens through selective social learning, where agents copy from agents with high payoffs [30, 32]. Motivated by empirical research on scientific [33] and technological [34] breakthroughs and previous studies on cultural evolution [5], we assume that high-quality (i.e. high-payoff) innovations are rare among many low-quality attempts to innovate. Hence, we draw the payoffs of newly innovated items, $z(x)$, from an exponential distribution, and subsequently square, double, and finally round values to an integer value [27]. This procedure results in a few traits with a payoff of around $z=50$, and roughly half of the traits with $z=0$ [26].
@@ 2.2 Dynamics
We consider a population of $N$ agents that are embedded in a social network. Initially, agents are naive, i.e. they do not possess any cultural traits, and their individual effort budget is $B$. A finite effort budget $B$ models the fact that individuals have limited resources, which also constrains cumulative cultural evolution at the collective level. In particular, over time agents spend their budget $B$ on actions to acquire cultural traits, either by copying already discovered traits (learning), or by discovering new traits (innovation). The dynamics of a focal agent $i$ at time $t$ (at development level $l_{i}$ and remaining resources $B(t)$) can be described as a cycle of the following three events:
@@
Agent $i$ selects its neighbour $j$ with the highest cumulative payoff [7]. If agent $j$ is at a higher development level than agent $i$, i.e. $l_{j}>l_{i}$, agent $i$ copies the $(l_{i}+1)$-stage trait of agent $j$, and the learning cost $C_{s}$ is subtracted from agent $i$’s budget. If the copying of a trait is not successful, either because the resources of agent $i$ are not sufficient ($B(t)<C_{s}$) or because the level of the highest trait of agent $j$ is too low, i.e. $l_{j}<l_{i+1}$, the learning costs $C_{s}$ are not deducted from the agent’s budget. Agent $i$ then continues with the innovation step (see below). If the social learning step was successful, the innovation step is skipped. Social learning is schematically depicted in Fig. 1C.
@@
If social learning was unsuccessful, and agent $i$ has a sufficient budget ($B(t)\geq C_{i}$), it will proceed with innovation. First, $i$ randomly selects a cultural branch $x$ and attempts to innovate a trait on level $l_{i}+1$. If the trait is viable, i.e. its payoff $z(x_{l_{i}+1})>0$, the innovation step is successful and agent $i$ acquires the trait. Note that unsuccessful innovation concludes the time step for agent $i$ without acquiring a new item. Regardless of the success of the innovation step, the cost of innovation $C_{i}$ is subtracted from agent $i$’s effort budget. Innovation is schematically depicted in Fig. 1B.
To quantify the performance of a population with respect to cultural accumulation, we calculate the mean payoff of all agents in the population, i.e. $\bar{Z}(t)=N^{-1}\sum Z_{i}(t)$, where $Z_{i}$ is the cumulative payoff of agent $i$, defined as $Z_{i}=\sum_{l=1}^{l_{\mathrm{max}}}z(x_{l})$. Note that to compute $Z_{i}$, we only sum over the complexity level $l$, as a consequence of the assumed dynamics of innovation and learning: an agent can only acquire a maximum of one trait per complexity level $l$. Additionally, we assume that the cost of acquiring a single trait depends only on the action performed, i.e., learning or innovation, but is independent of other factors, including the payoff $z$ and the complexity level $l$ of the trait. Furthermore, we follow the intuition that it is easier to copy something from a peer than to innovate it, i.e. we assume $C_{s}<C_{i}$ [26].
@@ 2.3 Algorithmic mediation
To investigate how algorithmic mediation can affect cultural accumulation, and to shed light on the interplay between different types of information propagation, i.e. social and algorithm-mediated learning, we extend the model described in the previous section. While still being embedded in a social network, agents can now also learn cultural traits through algorithmic mediation. For simplicity, we implement algorithmic mediation as a ’perfect’ recommender system that has full knowledge about all cultural traits innovated so far in the system, i.e. the cultural reservoirs of every agent. The knowledge about traits’ development levels $(l)$ and their payoffs ($z(x_{l})$) allows the algorithm to provide individuals with personalized recommendations. We assume that learning through algorithmic mediation requires less effort than social learning, i.e., we assume $C_{r}<C_{s}$, where $C_{r}$ defines the costs of algorithmic mediation. In each time step, the agent copies a trait through algorithmic mediation with probability $r$, or otherwise, with probability ($1-r$), the agent learns socially.
Figure 1: Schematic depiction of key model components. A: Open-ended cultural space depicted with four branches ($X=4$), and an infinite number of (complexity) levels $l$. B: Agents innovate a new item on a randomly chosen branch on a level $l+1$. C: Social learning is modeled as a two-step process. (1) Agents choose their neighbor $j$ with the highest cumulative payoff $Z_{j}$. (2) If agent $j$ possesses an item on a sufficient level ($l+1$), the item is copied by agent $i$ and the cost $C_{s}$ is paid. D: Algorithmic mediation is also modeled as a two-step process. (1) The recommender system aggregates all existing cultural traits and sorts traits on level ($l+1$) according to their payoffs $z$. (2) Agent $i$ then receives a recommendation of the trait with the highest payoff.
@@
In every time step, the recommender system aggregates all cultural traits and selects only those on level $l_{i}+1$ (personalization). After sorting those $l_{i}+1$-level traits in descending order (according to their payoffs) it recommends the highest-payoff trait, which is acquired by agent $i$ and the cost $C_{r}$ is deducted from its budget. Note that when no suitable traits can be recommended, or if agent $i$ does not have enough resources ($B(t)<C_{r}$), the recommendation fails, the agent does not pay a cost, and continues with the innovation step. When the recommendation is successful, agent $i$ does not attempt to innovate. Algorithmic mediation is schematically depicted in Fig. 1D.
@@
Agent $i$ attempts to copy a cultural trait from neighbors (the social learning step is identical to the one described in the previous section).
The order of events in each time step is identical to the one described in the previous section, i.e., consisting of $(i)$ the death-birth process, $(ii)$ learning (either socially or through algorithmic mediation), and $(iii)$ innovation. The different model components are schematically depicted in Fig. 1.

### v2

@@
Agnieszka Czaplicka1,2,∗, Fabian Baumann1,3, Iyad Rahwan1
@@
2Faculty of Physics, Warsaw University of Technology, Warszawa, Poland.
3Department of Biology, University of Pennsylvania, Philadelphia, PA 19104.
@@
The remarkable ecological success of humans is often attributed to our ability to develop complex cultural artifacts that enable us to cope with environmental challenges. The evolution of complex culture (cumulative cultural evolution) is usually modeled as a collective process in which individuals invent new artifacts (innovation) and copy information from others (social learning). This classic picture overlooks the growing role of intelligent algorithms in the digital age (e.g., search engines, recommender systems, large language models) in mediating information between humans, with potential consequences for cumulative cultural evolution. Building on a previous model, we investigate the combined effects of network-based social learning and a simplistic version of algorithmic mediation on cultural accumulation. We find that algorithmic mediation significantly impacts cultural accumulation and that this impact grows as social networks become less densely connected. Cultural accumulation is most effective when social learning and algorithmic mediation are combined, and the optimal ratio depends on the network’s density. This work is an initial step towards formalising the impact of intelligent algorithms on cumulative cultural evolution within an established framework. Models like ours provide insights into mechanisms of human-machine interaction in cultural contexts, guiding hypotheses for future experimental testing.
@@ 1 Introduction
Humans’ ability to create complex cultural artifacts such as languages, scientific theories, art, and technology (including physical artifacts) is often seen as one of the most important aspects of our success as a species [1, 2]. An important driving force behind this cultural accumulation is the ability of humans to learn socially, i.e., copy information from their peers [1, 3, 4, 5]. Social learning is a highly adaptive process, influenced by cognitive biases that can manifest as different social learning strategies ranging from copying successful individuals (payoff bias) [6, 7, 8], to frequent cultural artifacts (context bias) [9, 10], to specific information (content bias) [11]. At the population level, social learning is a necessary condition for cumulative cultural evolution (or cultural accumulation), a process that describes the emergence of increasingly complex cultural artifacts through a ratchet effect: By preserving important information over time—i.e., from generation to generation—social learning allows collectives to produce cultural artifacts that lie outside of the scope of individuals [12, 13]. Yet the cultural accumulation can also be limited by factors such as population size [13], the structure of the underlying social network [14], or the social learning strategies that humans use [15].
In the digital age, cultural accumulation is increasingly shaped by algorithmic mediation [16]. For the purposes of this paper, we broadly define algorithmic mediation as a collection of processes through which intelligent systems aggregate and redistribute information on an unprecedented scale, thereby impacting cultural accumulation. These algorithms include recommender systems that influence consumer behavior [17], content-ranking mechanisms on social media platforms [18], and search engines like Google [19]. More recently, large language models (LLMs) (e.g., GPT, BERT, Gemini, and others) have emerged as transformative tools, reshaping how we acquire knowledge and develop skills. Unlike earlier recommender systems, which primarily tackled e-commerce challenges by focusing on consumer preferences, LLMs are likely influencing a far broader range of domains that are central to cultural accumulation. These include generating ideas [20], facilitating problem-solving [21], assisting with technical tasks such as coding [22], and offering personalized explanations in fields like mathematics and science [23]. By synthesizing vast repositories of information and tailoring information to individual needs, algorithmic systems have indeed the potential to impact cumulative cultural evolution by enhancing innovation rates across domains, including technology and basic science [24], reinforcing filter bubbles [25], and limiting the diversity of information shared [26]. Recognizing these impacts, previous theoretical and empirical works on socio-technological systems have mainly focused on the impact of algorithms on the diversity of content [27], the formation of opinions [28, 29], and the spread of misinformation [30]. However, the question of how algorithmic mediation may affect cultural accumulation received much less attention.
@@
To close this gap, we build on—and extend—an established model of cultural evolution [31], where algorithmic mediation is conceptualized as a mechanism that aggregates and redistributes cultural traits on a collective level. By integrating social learning and algorithmic mediation in a single modeling framework, we are able to capture the dynamics induced by the interaction of localized peer influence (social learning) and global information redistribution (algorithmic mediation). Our results shed light on an emerging trade-off that humans face in the digital age: cultural accumulation is neither optimal through pure social learning nor through purely algorithmically mediated information. Instead, both types of learning are most powerful when combined, where the optimal ratio depends on the underlying social network. This insight underscores the need to examine algorithmic mediation and social learning as interconnected forces shaping cultural evolution in the digital age, rather than as isolated processes.
@@ 2 Model
We build on prior theoretical work that explored how cultural accumulation is influenced by various aspects of individual learning, including social learning strategies and the costs associated with learning [31]. First, we extend this model by incorporating an explicit social network structure. In a second step, we extend the model by a simple algorithm that mediates information exchange among agents.
The individual-based model presented in [31] simulates a dynamic process that unfolds over generations, where each ”time step” represents the simultaneous birth and death of an entire generation of $N$ agents. In this framework, agents first copy information from the previous generation and then, after acquiring all available knowledge, begin to innovate. In this work, we deviate from this approach and modify the model by re-scaling the dynamics: at each time step, a single cultural trait is acquired either through copying or innovation [32]. This modification enables us to examine the interplay between different mechanisms of information transmission—social learning and algorithmic mediation—and their implications for population-scale cultural accumulation.
Below, we describe our modeling approach and provide additional implementation details in the Supplemental Material (SM).
@@ 2.1 Cultural space
We consider a space of cultural traits (or cultural space) that defines how new traits (innovations) arise based on existing ones. In line with previous works, the cultural space consists of several cultural branches (or paths), and traits are functionally dependent, reflecting the cumulative nature of human culture [14, 33, 31, 34].
The cultural space is structured as follows. Each trait $x_{l}$ is defined by two components: the cultural branch $x$ it belongs to and its complexity level $l$, see Fig. 1A. Cultural branches $x$ model the fact that cultural evolution often follows multiple parallel paths [33, 14], and levels $l$ model the functional dependency and increasing complexity of cultural traits built on top of each other. We consider a finite set of possible branches, i.e., $x=1,...,X$ and assume that cultural complexity levels $l$ are unbounded, i.e. $l=1,2,...,\infty$, i.e. cultural artifacts can be modified infinite number of times. The functional dependency of the traits $x$ additionally requires that the traits are acquired in ascending order, starting with the lowest level $l=1$, or more generally: an agent that wants to acquire a trait on level $l$ must possess a trait on level ${l-1}$. Each cultural trait $x_{l}$ is characterised by an inherent quality score, or payoff, $z(x_{l})$, and we assume that the extent to which the traits are beneficial to agents increases with $z$ [5, 14, 35, 36]. Note, however, that we do not link an individual’s probability to survive to their payoff. Instead, cultural selection happens through selective social learning, where agents copy from agents with high payoffs [35, 37]. Motivated by empirical research on scientific [38] and technological [39] breakthroughs and previous studies on cultural evolution [5], we assume that high-quality (i.e. high-payoff) innovations are rare among many low-quality attempts to innovate. Hence, we draw the payoffs of newly innovated items, $z(x)$, from an exponential distribution, and subsequently square, double, and finally round values to an integer value [32]. This procedure results in a few traits with a payoff of around $z=50$, and roughly half of the traits with $z=0$ [31].
The distinction between complexity levels $(l)$ and payoffs $(z)$ in the model captures the relationship between the accumulation of culture (i.e. knowledge, skills, or innovations) and its practical utility in real-world cultural dynamics. For instance, consider the development of a new scientific method (skill) or technological tool (physical artifact). The complexity of a tool—reflected in its levels—might represent the number of iterative modifications and refinements over time, such as the progression from early calculators to modern computers. However, not all refinements enhance utility or payoff. Some modifications, like aesthetic improvements or added features, may increase the complexity of a tool without necessarily maximizing its functional effectiveness or societal impact. Similarly, in scientific endeavors, researchers may pursue increasingly complex methodologies that yield incremental theoretical insights, but only some of these contribute substantially to practical applications or yield high-impact results capable of shifting the paradigm within a discipline. This distinction between levels and payoffs underscores the importance to disentangle how cumulative cultural processes lead to either optimal or suboptimal outcomes, depending on the social and algorithmic mechanisms mediating cultural transmission.
@@ 2.2 Dynamics
We consider a population of $N$ agents that are embedded in a social network. Initially, agents are naive, i.e. they do not possess any cultural traits, and their individual effort budget is $B$. A finite effort budget $B$ models the fact that individuals have limited resources, which also constrains cumulative cultural evolution at the collective level. Over time, agents acquire cultural traits either by copying existing traits (learning) or by discovering new ones (innovation), with each action incurring a cost. The dynamics of a focal agent $i$ at time $t$, characterized by the highest complexity level $l_{i}$ of their traits and their remaining resources $B(t)$, can be modeled as a cycle consisting of the following three events:
@@
Agent $i$ selects their neighbour $j$ with the highest cumulative payoff [7]. If agent $j$ is at a higher complexity level $l$ than agent $i$, i.e. $l_{j}>l_{i}$, agent $i$ copies the $(l_{i}+1)$–stage trait of agent $j$, and the learning cost $C_{s}$ is subtracted from agent $i$’s budget. If copying is not successful, either because the resources of agent $i$ are not sufficient ($B(t)<C_{s}$) or because the level of the highest trait of agent $j$ is too low, i.e. $l_{j}<l_{i+1}$, the learning cost $C_{s}$ is not deducted from the agent’s budget. Agent $i$ then continues with the innovation step (see below). If the social learning step was successful, the innovation step is skipped. Social learning is schematically depicted in Fig. 1C. We mainly focus on payoff-biased learning as described above, however, in the SM we also show results for other social learning strategies.
@@
If social learning was unsuccessful, and agent $i$ has a sufficient budget ($B(t)\geq C_{i}$), they will proceed with innovation. First, agent $i$ randomly selects a cultural branch $x$ and attempts to innovate a trait on level $l_{i}+1$. If the trait is viable, i.e. its payoff $z(x_{l_{i}+1})>0$, the innovation step is successful and agent $i$ acquires the trait. Note that unsuccessful innovation concludes the time step for agent $i$ without acquiring a new item. Regardless of the success of the innovation step, the cost of innovation $C_{i}$ is subtracted from agent $i$’s effort budget. Innovation is schematically depicted in Fig. 1B.
To analyze cultural accumulation in the model, we use three distinct measures. First, we calculate the mean payoff of all agents in the population, defined as $\bar{Z}(t)=\frac{1}{N}\sum\limits_{i}Z_{i}(t)$, where $Z_{i}$ represents the cumulative payoff of agent $i$, given by $Z_{i}=\sum\limits_{l=1}^{l_{\mathrm{max}}}z(x_{l})$. Here, $z(x_{l})$ denotes the payoff associated with trait $x_{l}$, and the summation is over the complexity levels $l$, reflecting the model’s dynamics: each agent can acquire at most one trait per complexity level. The mean payoff $\bar{Z}(t)$ serves as a proxy for the performance of the population, indicating the overall utility derived from the accumulated traits. Second, we count the total number of distinct traits created during the process. This measure provides an indicator of the diversity of traits in cultural accumulation. Finally, we track the highest levels of traits reached by agents, which assess the average cultural complexity within the population. These three measures—performance, diversity, and complexity—together offer a comprehensive framework for analyzing the dynamics of cultural accumulation in the model.
Note that we assume that the cost of acquiring a single trait depends only on the action performed, i.e., learning or innovation, but is independent of other factors, including the payoff $z$ and the complexity level $l$ of the trait. Furthermore, we follow the intuition that it is easier to copy something from a peer than to innovate it, i.e. we assume $C_{s}<C_{i}$ [31].
@@ 2.3 Algorithmic mediation
To investigate how algorithmic mediation affects cultural accumulation and the interplay between social and algorithm-mediated learning, we extend the model described in the previous section. While still embedded in a social network, agents can now also learn cultural traits through algorithmic mediation. We implement this as a ’perfect’ algorithm with full knowledge of all cultural traits innovated so far, including their development levels $(l)$ and payoffs ($z$), enabling personalization of information. Learning through algorithmic mediation is assumed to be less costly than social learning ($C_{r}<C_{s}$). We implement the balance between social and algorithmic learning as a stochastic process: At each time step, an agent copies a trait via algorithmic mediation with probability $r$ or learns socially with probability $(1-r)$.
Figure 1: Schematic depiction of key model components. A: Open-ended cultural space depicted with four branches ($X=4$), and an infinite number of (complexity) levels $l$. B: Agents innovate a new item on a randomly chosen branch on a level $l+1$. C: Social learning is modeled as a two-step process. (1) Agents choose their neighbour $j$ with the highest cumulative payoff $Z_{j}$. (2) If agent $j$ possesses an item on a sufficient level ($l+1$), the item is copied by agent $i$ and the cost $C_{s}$ is paid. D: Algorithmic mediation is also modeled as a two-step process. (1) The algorithm aggregates all existing cultural traits and sorts traits on level ($l+1$) according to their payoffs $z$. (2) Agent $i$ then receives a recommendation of the trait with the highest payoff.
@@
In every time step, the algorithm aggregates all cultural traits and selects only those on level $l_{i}\!+\!1$ (personalization). After sorting those $l_{i}\!+\!1$–level traits in descending order (according to their payoffs) it recommends the highest-payoff trait, which is acquired by agent $i$ and the cost $C_{r}$ is deducted from their budget. Note that when no suitable traits can be recommended, or if agent $i$ does not have enough resources ($B(t)<C_{r}$), the recommendation fails, the agent does not pay a cost, and continues with the innovation step. When the recommendation is successful, agent $i$ does not attempt to innovate. Algorithmic mediation is schematically depicted in Fig. 1D.
@@
Agent $i$ attempts to copy a cultural trait from their neighbours, as described in the previous section.
The order of events in each time step is identical to the one described in the previous section, i.e., consisting of $(i)$ the death-birth process, $(ii)$ learning (either socially or through algorithmic mediation), and $(iii)$ innovation. The different model components are schematically depicted in Fig. 1. Details of the implementation are presented in the SM.
While idealized, our model of algorithmic mediation broadly mirrors the capabilities of advanced algorithmic systems, such as large language models (LLMs), which tailor information delivery to align with user needs in order to maximize utility, although their effectiveness remains constrained by current methodological and technological limitations. For instance, reinforcement learning from human feedback, as employed in systems like ChatGPT, enables the prioritization of useful, accurate, and contextually relevant information. While this process is not directly analogous to biological fitness proxies in traditional cultural evolution models, it effectively captures the adaptive refinement and redistribution of cultural traits—including knowledge and skills—within an algorithmically mediated environment.